OpenStudy (anonymous):

If f(x) = g(u) and u = u(x) then (A) f '(x) = g '(u) (B) f '(x) = g '(u) . u '(x) (C) f '(x) = u '(x)

OpenStudy (anonymous):

This is the chain rule. Do you know it?

OpenStudy (anonymous):

To be concrete, let's say f(x) = y = (1 + 2x)^2 let u = 1 + 2x y = u^2 How to get dy/dx?

OpenStudy (anonymous):

yeah.... :) The problem is that i dunno calculus.... its not even in my syllabus

OpenStudy (anonymous):

Do you know how to do derivatives by the power rule?

OpenStudy (anonymous):

Well, think of it this way: dy/dx = dy/du times du/dx we can *cancel* the du's

OpenStudy (anonymous):

okay...... then

OpenStudy (anonymous):

So for my example we do dy/du = 2u du/dx = 2 dy/dx = dy/du du/dx = 4u = 4(1 + 2x) which you can check by multiplying out.

OpenStudy (anonymous):

so is the answer (b) correct??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

owhkay...thank you so much !!

OpenStudy (anonymous):

yw